Solve the differential equation
d^2y/dx^2-2dy/dx+y=x^2e^2x
1
Expert's answer
2020-12-08T09:24:23-0500

For the given differential equation, A.E. is given by,

"D^2-2D+1=0"

Solving it, we get, "D=1,1"

Hence, "y=ae^{x}+bxe^{x}" (1)


P. I. "\frac{1}{D^2-2D+1}x^2e^{2x}=\frac{1}{(D-1)^2}x^2e^{2x}= e^{2x}\frac{1}{(D+2-1)^2}x^2"


"=e^{2x}\frac{1}{(D+1)^2}x^2= e^{2x}(D+1)^{-2}x^2"


"e^{2x}(1-2D+3D^3+.....)x^2 = e^{2x}(x^2-4x+6)" (2)

So total solution for the equation is,

"y=ae^{x}+bxe^{x}+e^{2x}(x^2-4x+6)"